OpenStudy (anonymous):

At a County fair there is a spinner game with 12 sectors. 2 red sector, 2 green sector, 2 blue sector, and 6 yellow sectors. If the spinner lands on a red sector, the player wins 2 tokens. If the spinner lands on a green sector, the player wins 2 tokens. If the spinner lands on a blue sector, the player wins 2 tokens. If the spinner lands on a yellow sector, the player loses 3 tokens. Is this game fair for the player and how much will the player win or lose on an average over time?

OpenStudy (kropot72):

Let the token profit be X. For a win, the profit is 2 tokens. For a loss, the profit is -3 tokens. The probability of a win is \[\frac{2}{12}+\frac{2}{12}+\frac{2}{12}=\frac{6}{12}=\frac{1}{2}\] the probability of a loss is \[\frac{6}{12}=\frac{1}{2}\] Setting up a probability distribution variable gives: \[E(X)=(-3\times \frac{1}{2})+(2\times \frac{1}{2})=-1\frac{1}{2}+1=-\frac{1}{2}\ token\] Therefore, on average, the player would expect to lose a half a token per game.